OpenStudy (anonymous):

A weak acid HA (Ka=1.0x10-4) is titrated with NaOH. The concentration of NaA at the equivalence point is 0.0001 M. What is the pH at equivalence point?

OpenStudy (vincent-lyon.fr):

So, you have to find out the pH of a weak base solution of concentration 10-4 M, whose pKa = 4 (or pKb = 10)

OpenStudy (anonymous):

Can you please explain further?

OpenStudy (vincent-lyon.fr):

Do you know how to work out the pH of a weak acid or weak base? If not, you cannot solve this problem. Have a look at your textbook.

OpenStudy (anonymous):

Kb= 1.0 x 10^-14/ 1.0 x 1.0 x 10^-4 = 1.0 x 10^-10 Now [OH-] = \[\sqrt{1.0 \times 10^{-10} \times 0.0001}= 1.0 \times 10^{-7}\]

OpenStudy (anonymous):

pH= 14-(-log1.0x10^-7)=7

OpenStudy (anonymous):

Did I do it correctly? :O

OpenStudy (vincent-lyon.fr):

Yes, you did: the base is so weak and the concentration so small that your pH is that of (almost) pure water.

OpenStudy (vincent-lyon.fr):

If you need full derivation for this case, I can help tomorrow.

OpenStudy (anonymous):

Yes, sure. And thank you so much! :)